

********************************************************************
UARS WEEKLY STATUS REPORT - 02 JAN 2004
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate in good health gathering solar data on
all available orbital tracking periods.
***********
SUSIM
***********
SUSIM successfully gathered all of its daily mid (1.1 nm)
resolution, Mg II, and high resolution monitoring scans
during this past week for periods of UARS real-time telemetry
or onboard tape recording.
***********
PEM
***********
Level 2 PEM HEPSA data files are being reprocessed for submission to the
DAAC. This data has been designated as version 2; however, there is no format
difference from the version 1 data format. Version 2 HEPSA data is being
generated because it was discovered that there exists processing errors within
some of the HEPSA version 1 data files. Thus, if you are using any HEPSA
version 1 data files from the DAAC, please replace them with the version 2
data files. Currently, HEPSA data between 1991274 and 2003119 has been
reprocessed and archived at the GSFC DAAC.
Level 2 PEM MEPS data file documentation is complete and submitted to
the DAAC. MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 1992209.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
2003 was a successful operational year for HALOE. The following
instrument performance values continue to be as expected for 2003:
- Science channel end to end trends of solar exoatmospheric values
- Instrument temperatures and voltages values
- gain and boresight test results
Use of the tape recorder to capture events, that do not occur during
TDRS contacts, provides the capability to capture most of the sunrise
and sunset events except during few days during the year. The third
TDRS will hopefully fill in the coverage in addition to provide more
opportunities for other instruments. The third TDRS may also provide
additional power management options during low beta angles by limiting
use of the transmitter during solar array night.
Science data processing at LaRC continues successfully, but data
contamination, primarily from zero fill, provides processing and data
certification challenges. However the challenges are worthwhile. The
science data products continue to be excellent and many instrument teams
use the HALOE data for comparisons.
During a time of limited resources the UARS Flight Operations Team has
performed superbly. In addition to keeping an aging spacecraft fling
true, they have provided the extra effort required now to support
instrument operations. Tape recorder scheduling has been near flawless
and the many power-on's has been extremely smooth. Our thanks to you
from the HALOE Operations Team . Happy New Year.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 20 December 2003 (Orbit 67150, GMT Day 354) through
26 December 2003 (Orbit 67254, GMT Day 360).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 356/1233 67187 AXIS On
PEM 356/1419 67188 AXIS Off
PEM 356/1419 67188 AXIS On
HALOE 357/0253 67196 Power Off
PEM 359/1308 67232 Fulltime
PEM 360/2145 67253 AXIS-2 Off
The beta angle ranged between the angles of 33.2 degrees to 63.4
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 80 degrees on 01 January 2004 (DOY 001).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.63 to 0.94 degrees C. Additional
battery performance data for each battery on 19 December and 26
December 2003 are provided in Appendix D.
The clock error ranged from 10.5 to -10.1 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
PEM = 05
FOT = 03
HALOE = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 00
Five (5) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
No operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The three new desktops have been created at JPL to support the three
possible configurations needed for UARS DSN events. Now that these
desktops have been made, another spacecraft test has been scheduled
for 30 December 2003.
The FOT continued to investigate the large attitude errors seen on the
simulator during each run of the day-in-the-life (DITL). The power
issues that caused the attitude to diverge were corrected but
subsequent runs still showed the spacecraft (on the simulator) lose
control. More analysis is needed to determine the reason behind the
large errors. This problem needs to be resolved so that the DITL
testing can continue.
MEETINGS
-None this report period
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 20 January 2004 (DOY 020).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 26 December 2003
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
67183 356/0631 03-329 OBC CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
67187 356/1233 03-330 PEM AXIS POWER ON
67188 356/1419 03-331 PEM AXIS POWER OFF
67188 356/1419 03-330 PEM AXIS POWER ON
67196 357/0235 03-332 HALOE POWER OFF
67206 357/1930 03-333 OBC CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
67232 359/1308 03-334 PEM DAY-ONLY TO FULLTIME
67235 359/1741 03-335 OBC CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
67253 360/2145 03-336 PEM AXIS-2 POWER OFF
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 26 December 2003
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 26 December 2003
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
67183 356/0631 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
67206 357/1930 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
67235 359/1741 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 26 December 2003
19 December 2003 - GMT Day 353 Beta = 33.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.31 +00.0/+00.0 07.2 0.00 0.0
2 +11.2/-16.8 +5.46/+4.50 +28.4/-13.6 25.0 1.03 20.6
3 +16.8/-22.4 +4.50/+3.87 +27.6/-13.6 24.8 1.02 20.3
26 December 2003 - GMT Day 360 Beta = 63.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.67/-4.62 +00.0/+00.0 07.2 0.00 0.0
2 +11.2/-22.4 +3.87/+2.91 +20.4/-15.2 26.4 1.02 17.5
3 +11.2/-33.6 +2.91/+1.97 +20.0/-14.8 26.4 1.01 17.0
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period
APPENDIX E
UARS ANOMALY REPORTS
Week ending 26 December 2003
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
03-055 FOT See List Dec 03 30832 Generic Late Acquisition
03-056 FOT 66992 343/1259 34488 Data Drops
03-057 ATC 67044 347/0100 N/A S/W Stored Cmd Error
03-058 HRDI 67059 348/0101 N/A Not accepting loads
03-059 FOT 67098 350/1515 34541 TDRSS East MA Problem